OpenStudy (anonymous):

Find the rectangular coordinates of the polar coordinate. (2/3, -2pi/3)

OpenStudy (anonymous):

\[x=r*cos(\theta)\\y=r*sin(\theta)\]We have r and theta\[x=\frac{2}{3}*cos\left(-\frac{2\pi}{3}\right)\\y=\frac{2}{3}*sin\left(-\frac{2\pi}{3}\right)\]And our Rectangular coordinate is\[\left(-\frac{1}{3},-\frac{\sqrt{3}}{3}\right)\]